OpenStudy (anonymous):

if 7x+2y-6z = 12 and if x,y and z are positive, then what is the value of 2+z/ 7x+2y? A) 1/12 B) 1/6 C) 1/4 D) 5/12 E) 7/12

OpenStudy (amistre64):

hitting the enter key helps to make this more readable to those of us who want to help ...

OpenStudy (anonymous):

sorry... I did, but it turned out like that

OpenStudy (amistre64):

:) fair enough

OpenStudy (amistre64):

if 7x+2y-6z = 12 ;x,y and z are positive what is the value of (2+z)/(7x+2y) is the question right? A) 1/12 B) 1/6 C) 1/4 D) 5/12 E) 7/12

OpenStudy (anonymous):

yeah, thanks

OpenStudy (id21):

I can't find the answer. Only (2+z)/6z

OpenStudy (amistre64):

i beleive we have to construct a plane from this ....

OpenStudy (amistre64):

7x+2y-6z -12 = 0 trying to recall my vector stuff :)

OpenStudy (amistre64):

7(x-Px) +2(y-Py) -6(z-Pz) = 0 such that ..... -7Px -2Py +6Pz = -12

OpenStudy (amistre64):

7x+2y-6z = 12 ; /-6 and solve for z to get -7x/6 -y/3 +z = -2 f(x,y) = z = 7x/6 +y/3 -2 gf = ( 7/6 , 1/3 )

OpenStudy (amistre64):

got know idea but this might work; sub in 7/6 for x and 1/3 for y z = 7(7)/6(6) +(1)/3(3) -2 z = 49/36 +1(4)/9(4) -2(36)/(36) z = (49 +4 -72)/36 = -19/36 ....... guess that is the wrong route to take :)

OpenStudy (anonymous):

... o.o"

OpenStudy (amistre64):

it might help to know what the section this is in is talking about ...

OpenStudy (anonymous):

... I know its an SAT question... but asides from that, I have no idea how to do it, or anything about it....